OpenStudy (anonymous):

verify using trig identities: 3sinx-2cos^2x/2+sinx=2sinx-1

OpenStudy (raden):

use the identity : cos^2(x) = 1 - sin^2(x)

OpenStudy (raden):

so, for the numerator can be 3sin(x) - 2cos^2(x) = 3sin(x) - 2(1- sin^2(x)) = 2sin^2(x) + 3sin(x) - 2

OpenStudy (raden):

factor out that, then u will can simplify by do cancel out for the similar terms

OpenStudy (raden):

got it ?

OpenStudy (anonymous):

Yeah thanks for the help.